OpenStudy (anonymous):

A person standing close to the edge on the top of a 160-foot building throws a baseball vertically upward. the quadratic function s(t) = -16t^2+64t+160 models the ball's height above the ground, s(t), in feet, t seconds after it was thrown a) after how many seconds does the ball reach its maximum height? What is the maximum height? b) How many seconds does it take until the ball finally hits the ground? Round to the nearest tenth of a second so I have found a) as 2 seconds and 224 feet but how do i figure out part b?

OpenStudy (jim766):

axis of symetry -b/2a -64/-32 = 2 x = 2 is a vertical line that goes through the max in this case x = 2 seconds when it reaches max height questions?

OpenStudy (anonymous):

I got that part I just need for part b saying how many seconds does it take until the ball finally hits the ground ... round to the nearest 10th

OpenStudy (jim766):

you can use two methods, you could graph it on a graphing calculator and see what x =, when the graph crosses the x axis. Doing it by hand, When the rock hits the ground the height will be zero, so set the equation equal to zero. 0 = -16t^2+64t+160 now factor it 0 = -16(t^2 - 4t -10) you would then use the quadratic formula to factor t^2-4t-10 Honestly, it is much easier to use the calculator and graph it if you can.